OpenStudy (anonymous):

find the area between y=x/(x^(2)+1), the x axis, x=1 and x=10

OpenStudy (anonymous):

jst do intgration

OpenStudy (anonymous):

342 is ans

OpenStudy (anonymous):

ya its integration, but i don't understand the x axis part. is it y=0 or something?

OpenStudy (anonymous):

1.96 is the correct ans my mistake

OpenStudy (zarkon):

\[\int_{1}^{10}\frac{x}{x^2+1}dx\] let \[u=x^2+1\]

OpenStudy (anonymous):

so does the x axis part not really matter? it's just a boundary line and i just need to use the x=10- and x=1?

OpenStudy (anonymous):

integration with dx itself means it is with x axis

OpenStudy (zarkon):

you could also write it like this \[\int_{1}^{10}\left(\frac{x}{x^2+1}-0\right)dx\]

OpenStudy (zarkon):

\[\int_{a}^{b}\left(top-bottom\right)\,dx\]

OpenStudy (anonymous):

ooo...now it clicks. thanks!!!